# AweDist: Attention-aware Embedding Distillation for New Input Token Embeddings

### v1

@@ Abstract
Current language models rely on static vocabularies determined at pretraining time, which can lead to decreased performance and increased computational cost for domains underrepresented in the original vocabulary. New tokens can be added to solve this problem, when coupled with a good initialization for their new embeddings. However, existing embedding initialization methods either require expensive further training or pretraining of additional modules. In this paper, we propose AweDist and show that by distilling representations obtained using the original tokenization, we can quickly learn high-quality input embeddings for new tokens. Experimental results with a wide range of open-weight models show that AweDist is able to outperform even strong baselines111Our code is available at https://github.com/konstantinjdobler/awedist..
@@ 1 Introduction
Pretrained language models are built upon static vocabularies that often fail to represent novel or domain-specific terminology, resulting in broken tokens, longer input sequences, and diluted semantic representations. This excessive tokenization not only leads to reduced performance on downstream tasks (Rust et al., 2021) but also increases the computational (and therefore also financial) cost (Ahia et al., 2023). While adding new tokens to a model’s vocabulary can reduce over-tokenization, it is crucial to choose a good initialization for the new embeddings (Gee et al., 2022; Dobler and de Melo, 2023; Yamaguchi et al., 2024a).
In this paper, we argue that many recent methods for embedding initialization are fundamentally limited. Whenever we wish to add a new token to a pretrained model’s vocabulary, this new token may be split up into multiple subtokens in the original model’s tokenization. However, these subtokens might not be individually informative about the semantics of the entire new token (consider e.g. <_pal> <at> <able>). The semantic processing of these subtokens as a sequence will largely not be represented in their input embeddings at all – but rather in the attention mechanism and feed-forward network layer weights of the Transformer (Elhage et al., 2022; Lad et al., 2024).
Therefore, methods not exploiting information encoded in Transformer layer weights are at a serious disadvantage. Motivated by this insight, we propose a novel method for input embedding initialization that captures information stored in all Transformer layers in addition to the existing input embeddings. Our method optimizes new token embeddings with a distillation-based objective to match the model’s behavior compared to when a new token is split up into its original subtokens. We demonstrate the efficacy of our method, dubbed AweDist for Attention-aware Embedding Distillation, in Section 5. We illustrate our method in Figure 1 and describe it in Section 3.
We compare against strong baselines, including standard embedding initialization procedures, training of the embedding matrices with causal language modeling, as well as pretrained hyper-networks that predict new token embeddings. Extensive experiments on a wide range of open-weight models including question-answering benchmarks and the generation of definitions for the newly added tokens confirm the effectiveness of our approach. Our experimental setup is detailed in Section 4. Additionally, we describe related work in Section 2 and explicitly discuss limitations of our method in Appendix A. In summary, our contributions are:
We propose AweDist, a novel method for providing high-quality input embeddings when adding new tokens to pretrained language models.
We motivate our proposed method by describing fundamental limitations of current embedding initialization methods and empirically verify our claims.
@@ 2 Background
Most state-of-the-art Large Language Models (LLMs) are trained using a static tokenizer, usually derived by a byte-pair encoding scheme before model training (Sennrich et al., 2016). However, a single tokenizer predetermined before model training might not be equally well-suited for the many different use cases that a model will need to address later on. Suboptimal tokenization of new languages or domain-specific text increases the cost of further training (Ahia et al., 2023; Yamaguchi et al., 2024a) and has also been tied to reduced performance (Rust et al., 2021; Ali et al., 2024).
@@
A solution to this problem is to modify the existing vocabulary to suit the specific needs. However, simply modifying the vocabulary is insufficient – We also need suitable embeddings for the new tokens. A common approach is simple random initialization followed by a phase of training only the new embedding layers on the target corpus to “warm up” the new embedding weights. However, random initialization of new token embeddings has been widely shown to underperform better methods (e.g., Gee et al., 2022; Minixhofer et al., 2022; Dobler and de Melo, 2023). This has led to notable interest in devising ways of obtaining more informative embeddings for new tokens.
@@
A common alternative is to initialize a new token’s embedding as the mean of its subtoken embeddings (Sachidananda et al., 2021; Koto et al., 2021; Gee et al., 2022) – potentially weighted based on the subtoken position in the new token (Nakash et al., 2025) – or based on other heuristics (Downey et al., 2023). Alternatively, a weighted mean of existing embeddings can also be computed based on similarity scores in auxiliary embedding spaces (Wang et al., 2019; Tran, 2020; Minixhofer et al., 2022; Ostendorff and Rehm, 2023; Dobler and de Melo, 2023; Liu et al., 2024, inter alia). In fact, most recent research into embedding initialization for new tokens proposes some variation of a weighted mean or copy of existing embedding vectors (Mosin et al., 2023; Zeng et al., 2023; Mundra et al., 2024; Yamaguchi et al., 2024a, b; Remy et al., 2024). Nonetheless, the resulting embeddings still require further tuning to achieve good results (Minixhofer et al., 2024). We refer to this category of initialization methods as “model-gradient free”, as they do not consider any training signal from the model.222While auxiliary embedding spaces can be trained, they do not backpropagate through the target model.
@@
Only a few methods based on access to model gradients have been investigated in recent years. Model gradients can be used by directly optimizing just the new token embeddings to minimize a cross-entropy loss on relevant contexts (Lampinen and McClelland, 2018), the common practice of training all embeddings while freezing other parts of the model (Artetxe et al., 2020; de Vries and Nissim, 2021) – potentially only backpropagating through parts of the model (Marchisio et al., 2023) – or via (pre-)training a hyper-network (Ha et al., 2017) that is able to predict embeddings for new tokens in the target embedding space (Pinter et al., 2017; Schick and Schütze, 2019, 2020; Teehan et al., 2024; Minixhofer et al., 2024). These methods – hyper-networks in particular – have been shown to yield immediately useful representations but require additional compute for (pre-)training. In the case of hyper-networks, it additionally might be necessary to train different hyper-networks for different target domains, further increasing the complexity.333Minixhofer et al. (2024) train separate hyper-networks for mistralai/Mistral-7b-v0.1 for English+Code and multilingual target tokenizers. Motivated by these difficulties, we propose an alternative method in this paper that does not require extensive pretraining of an embedding prediction hyper-network but still achieves competitive results.
3 Method: AweDist – Attention-aware Embedding Distillation
@@
3.1 Motivation
As illustrated in Section 1, individual subtoken embeddings (e.g., <_pal> <at> <able>) of a new token (e.g., <_palatable>) do not necessarily store the semantics of the specific concatenated form . Instead, their contexualized representation is gradually constructed in the Transformer layers via a neural detokenization process (Elhage et al., 2022; Lad et al., 2024) and then attended to by other token positions in the sequence, which are influenced by this contextualized representation. The various prior approaches that solely rely on information encoded in the embedding tables of a model thus ignore the crucial functional knowledge stored in the attention and feed-forward weights of Transformer layers, which imposes a fundamental limitation on their performance.
@@ 3.2 Method Description
Let be the tokenizer of a pretrained Transformer. Given a new token string with original subtokens , we require a small corpus of example sequences each containing the new token string . We denote the new tokenizer, which includes , as , and the tokenization of using or as and , respectively. The target for learning the embedding of are the hidden states at a specified layer produced by the language model when processing , which we denote as . We wish to find an embedding for a new token such that hidden states produced by processing the input sequence instead of are similar to our target . We denote these new hidden states using our new embedding as . As the two input sequence indices are not aligned due to replacing subtokens by a single new token, we define a set of mapped positions , where and are the positions of the same token in and , respectively. Additionally, only includes pairs where position in would attend to . We use subscripts such as to signify the hidden state at the -th token position. Therefore, our objective for a given target layer is defined as:
@@ 3.3 Further Details
Like other methods optimizing embeddings based on contexts, our method also needs input sequences. Firstly, we note that randomly sampling texts from a domain-specific or general corpus is inefficient for our goal of learning newly added embeddings: typically, the new tokens will only make up a small fraction, so most gradient updates will actually not affect the new input embeddings at all and simply learn to minimize their log-probability in the case of output embeddings. As we aim to have a fast method, we need a better approach. We propose two different approaches: (1) Our main approach is to simply retrieve snippets that contain our target tokens from a domain-specific or general corpus. This can be implemented efficiently using the algorithm proposed by Aho and Corasick (1975). Then we can truncate the snippets to a small window around our target token to optimize computational efficiency. (2) For causal language models, we can simply generate snippets by prompting the model with our target token. We provide implementation details in Section B.5. In our main experiments, we focus on the first approach but note the availability of the second approach in case a reference corpus is not available. In most cases of domain and language adaptation, the availability of such a corpus is a reasonable assumption. Nevertheless, we study an ablation instead using the generative approach in Section 5.2.
@@
Since our method backpropagates gradients back from the hidden states, we do not learn output embeddings with our distillation-based objective. In fact, this is not possible, as our new tokens are not part of the original model that serves as the “teacher”. In practice, for learning output embeddings, we can simply add a next-token prediction objective just for the output embeddings at a minimal computational overhead or combine our method with any other method for initializing output embeddings. Recent work suggests that input and output embeddings should in fact be treated differently (Nakash et al., 2025; Huang et al., 2025). Indeed, by no means does the Transformer architecture necessitate a shared vocabulary between the input and output embeddings. This is merely a convention that makes some use cases easier to implement. Therefore, we can either only add new tokens to the input embedding matrix or – for compatibility with existing frameworks – set their output embeddings to a vector of zeros.
@@
We employ a simple setup and use the AdamW (Kingma and Ba, 2017; Loshchilov and Hutter, 2019) optimizer for all trainable parameters. We set the batch size to optimize throughput and run all experiments on Nvidia H100 80GB GPUs. We do not use weight decay and maintain a constant learning rate with a linear warmup. For fair comparison, we sweep for the best learning rate for all methods that require a learning rate. Since our method is aimed to serve as an “initialization” rather than as full-scale further training, we restrict the number of example sequences to a maximum of 25 per target token and truncate to a context length of 50 tokens. These restrictions ensure that our method is quick to run, initializing 2,500 new tokens on a single GPU in under 10 minutes. We use the same data for all training-based methods, including baselines and variations of our method. We provide all details in Appendix B.
@@ 4.1 Evaluation
A common use case for adding new tokens to a pretrained model’s tokenizer is domain adaptation (Gururangan et al., 2020), especially for the biomedical domain (Poerner et al., 2020; Gee et al., 2023; Hasan et al., 2024; Singh et al., 2024). This is because it is a particularly challenging domain with highly complex domain-specific terminology, which can serve as a benchmark to stress test embedding initialization methods. Therefore, we evaluate our method on a collection of standard benchmarks in the biomedical domain (Pal et al., 2024) and add frequently occurring words as new tokens.
Additionally, for a more in-depth evaluation of the quality of new representations provided by our methods, we prompt the adapted models to generate definitions for new tokens and evaluate their quality with an LLM-as-a-Judge (Li et al., 2023). Furthermore, to push the limits of embedding initialization methods, we move beyond domain adaptation and apply our method to multi-word tokens, which have recently gained further interest (Gee et al., 2023; Liu et al., 2025; Huang et al., 2025). We follow the zero-shot tokenizer transfer setting in our evaluations (Minixhofer et al., 2024), i.e., we evaluate initialization methods without further training on a corpus, as we want to directly judge the quality of the resulting representations.
We select target tokens that actually occur frequently in the chosen benchmarks so that we can effectively judge the quality of newly generated embeddings. We select words that occur more frequently than a threshold and additionally ensure that all individual benchmarks are represented. We report the full methodology in Section B.1. For our multi-word token experiments, we prompt gpt-o4-mini-high to generate suitable candidates (full details in Section B.2). For our experiments in the biomedical domain, we retrieve contexts from ncbi/pubmed as a reference corpus. For our multi-word token experiments, we prompt the original models to generate sequences containing the new tokens. For the LLM-as-a-Judge evaluations, we use Llama-3.3-70B-Instruct as the judge model. We evaluate the general correctness of the generated definitions to test the quality and completeness of the resulting representations as well as semantic similarity with the target model’s output (Villegas et al., 2025), as inducing as little behavior change as possible can be an important desideratum.
@@ 4.2 Modeling
We conduct experiments using a wide range of open-weight model checkpoints to ensure our results are not merely a function of peculiarities in any specific model’s embedding space. In particular, we choose the following models: Mistral-7B-v0.1, OLMo-2-7B-1124-Instruct, Llama-3-8B, Llama-3-8B-Instruct, Llama-3.1-8B, Llama-3.1-8B-Instruct, Llama-3.2-3B, and Llama-3.2-3B-Instruct. In the remainder of the paper, we denote -Instruct variants with a simple -i. Through this extensive evaluation, we study different model families, model sizes (3B, 7B, and 8B), instruction-tuned and base models, as well as models with separate input/output embeddings and tied (shared) embeddings between input and output.
@@
For all results, we report the performance of the unmodified base model using the original tokenization, which serves as our “Target”. To establish a lower bound, we also report results for initializing the new token embeddings randomly from a normal distribution with a per-channel mean and standard deviation of the original embeddings (Hewitt, 2021). Furthermore, we report results for the commonly used method of taking the subtoken mean (Sachidananda et al., 2021; Koto et al., 2021; Gee et al., 2022) as initialization for a new token, which has been shown to perform similarly to more sophisticated initialization methods (Minixhofer et al., 2024) that also use a weighted average of existing embeddings.
Since our proposed method conducts a short optimization on a few reference sequences per new token, we compare against the common approach of training embeddings using causal language modeling using the same data. Specifically, we report results for “classic” embedding tuning using the next-token prediction objective (NTP) as well as masking updates to the original embeddings such that only new embeddings are optimized (NTP++), which corresponds to the method used by Lampinen and McClelland (2018). These methods use the subtoken mean as their starting point. A strong alternative method for obtaining embeddings for new tokens are hyper-networks specifically pretrained for this task. We report results for ZeTT (Minixhofer et al., 2024) using their provided hyper-network checkpoints. We run five random seeds for all methods that include randomness and report the mean and standard deviation.
@@
As the initial embedding input into our method, we also use the subtoken mean. We investigate alternatives to this choice in Section 5.2. Our proposed objective can also be combined with the NTP-based objectives. We report results for combining AweDist and NTP++. Since the NTP++ and AweDist objectives have different scales, a simple addition is suboptimal. In fact, NTP++ is usually of larger magnitude, which leads it to overpower AweDist. Therefore, we also consider an “autoscaled” variant AweDist-NTP++ where the NTP loss is scaled by = stop_gradient(loss_awedist / loss_ntp) before summing. Here, stop_gradient prevents gradient flow through the scaling factor .444Multi-objective optimization is a rich field with extensive literature, we leave a thorough investigation of alternatives such as GradNorm (Chen et al., 2018) for future work.

### v2

# Token Distillation: Attention-Aware Input Embeddings for New Tokens

@@ Abstract
Current language models rely on static vocabularies determined at pretraining time, which can lead to decreased performance and increased computational cost for domains underrepresented in the original vocabulary. New tokens can be added to solve this problem, when coupled with a good initialization for their new embeddings. However, existing embedding initialization methods require expensive further training or pretraining of additional modules. In this paper, we propose Token Distillation and show that by distilling representations obtained using the original tokenization, we can quickly learn high-quality input embeddings for new tokens. Experimental results with a wide range of open-weight models show that Token Distillation outperforms even strong baselines.111Our code is available at https://github.com/konstantinjdobler/token-distillation.
@@ 1 Introduction
Pretrained language models are trained with a fixed tokenizer that often fragments domain-specific or novel terms into multiple subtokens. This excessive tokenization not only leads to reduced performance on downstream tasks (Rust et al., 2021; Ali et al., 2024) but also increases the computational (and therefore also financial) cost due to inflated sequence lengths (Ahia et al., 2023; Yamaguchi et al., 2024a). Although adding new tokens to a model’s vocabulary can reduce over-tokenization, it is crucial to choose a good initialization for the new embeddings (Gee et al., 2022; Minixhofer et al., 2022; Dobler & de Melo, 2023; Yamaguchi et al., 2024a).
In this paper, we argue that many recent methods for embedding initialization are fundamentally limited. Whenever we wish to add a new token to a pretrained model’s vocabulary, this new token may be split up into multiple subtokens in the original model’s tokenization. However, these subtokens might not be individually informative about the semantics of the entire new token (consider, e.g., <_pal> <at> <able>). The semantics of a word composed of multiple subtokens will largely not be stored in their raw input embeddings at all – but rather constructed by the Transformer’s attention/feed-forward layer stack during contextualization (Elhage et al., 2022; Lad et al., 2024; Kaplan et al., 2025).
Therefore, methods that do not exploit the information encoded in Transformer layer weights are at a serious disadvantage. Motivated by this insight, we propose a novel method for input embedding initialization that captures information stored in all Transformer layers in addition to the existing input embeddings. Our method optimizes new token embeddings with a distillation-based objective to match the model’s behavior compared to when a new token is split up into its original subtokens. We demonstrate the efficacy of our method, dubbed “Token Distillation”, in Section 5. We illustrate our method in Figure 1 and describe it in Section 3.
We compare against strong baselines, including standard embedding initialization procedures, training of the embedding matrices with causal language modeling, as well as pretrained hyper-networks that predict new token embeddings. Extensive experiments on a wide range of open-weight models, including question-answering benchmarks and the generation of definitions for the newly added tokens confirm the effectiveness of our approach. Our experimental setup is detailed in Section 4. Additionally, we describe related work in Section 2 and explicitly discuss limitations of our method in Limitations. In summary, our contributions are as follows.
We propose Token Distillation, a novel method for providing high-quality input embeddings when adding new tokens to pretrained language models.
We motivate our proposed method by describing the fundamental limitations of current embedding initialization methods and empirically verify our claims.
@@ 2 Background
Most state-of-the-art Large Language Models (LLMs) are trained using a static tokenizer, usually derived by a byte-pair encoding scheme before model training (Sennrich et al., 2016). However, a single tokenizer predetermined before model training might not be equally well-suited for the many different use cases that a model will need to address later on. Suboptimal tokenization of new languages or domain-specific text increases the cost of further training and inference (Ahia et al., 2023; Yamaguchi et al., 2024a) and has also been tied to reduced performance (Rust et al., 2021; Ali et al., 2024). Furthermore, Lesci et al. (2025) show that in practice, words which are not a single token in the vocabulary are significantly less likely to be generated.
@@
A solution to this problem is to modify the existing vocabulary to suit the specific needs. However, simply modifying the vocabulary is insufficient – we also need suitable embeddings for the new tokens. A common approach is simple random initialization followed by a phase of training only the new embedding layers on the target corpus to “warm up” the new embedding weights. However, random initialization of new token embeddings has been widely shown to underperform better methods (e.g., Gee et al., 2022; Minixhofer et al., 2022; Dobler & de Melo, 2023). This has led to notable interest in devising ways of obtaining more informative embeddings for new tokens.
@@
A common alternative is to initialize a new token’s embedding as the mean of its subtoken embeddings (Sachidananda et al., 2021; Koto et al., 2021; Gee et al., 2022) – potentially weighted based on the subtoken position in the new token (Nakash et al., 2025) – or based on other heuristics (Downey et al., 2023). Alternatively, a weighted mean of existing embeddings can also be computed based on similarity scores in auxiliary embedding spaces (Wang et al., 2019; Tran, 2020; Minixhofer et al., 2022; Ostendorff & Rehm, 2023; Dobler & de Melo, 2023; Liu et al., 2024, inter alia). In fact, most recent research into embedding initialization for new tokens proposes some variation of a weighted linear combination or copy of existing embedding vectors (Mosin et al., 2023; Zeng et al., 2023; Mundra et al., 2024; Yamaguchi et al., 2024a; b; Remy et al., 2024; Goddard & Neto, 2025; Lee et al., 2025; Singh et al., 2025). Nonetheless, the resulting embeddings still require further tuning to achieve good results (Minixhofer et al., 2024). We refer to this category of initialization methods as “model-gradient free”, as they do not consider any training signal from the model.
@@
Only a few methods based on access to model gradients have been investigated in recent years. Model gradients can be used by directly optimizing just the new token embeddings to minimize a cross-entropy loss on relevant contexts (Lampinen & McClelland, 2018), the common practice of training all embeddings while freezing other parts of the model (Artetxe et al., 2020; de Vries & Nissim, 2021) – potentially only backpropagating through parts of the model (Marchisio et al., 2023) – or via (pre-)training a hyper-network (Ha et al., 2017) that is able to predict embeddings for new tokens in the target embedding space (Pinter et al., 2017; Schick & Schütze, 2019; 2020; Teehan et al., 2024; Minixhofer et al., 2024). Alternatively, using PatchScopes (Ghandeharioun et al., 2024), “Tokens to Words” (Kaplan et al., 2025) locates layers at which a word fragmented into multiple subtokens is represented by a single contextual embedding in the residual stream, and then projects these hidden states into the input/output embedding spaces via trained mappings. These methods – hyper-networks in particular – have been shown to yield immediately useful representations but require additional compute for (pre-)training. In the case of hyper-networks, it additionally might be necessary to train different hyper-networks for different target domains, further increasing the complexity.222Minixhofer et al. (2024) train separate hyper-networks for mistralai/Mistral-7b-v0.1 for English+Code and multilingual target tokenizers. Motivated by these difficulties, we propose an alternative method in this paper that does not require extensive pretraining of an embedding prediction hyper-network or additional modules and still achieves competitive results, even outperforming pretrained hyper-networks.
3 Method: Token Distillation
@@
3.1 Intuition
As illustrated in Section 1, individual embeddings for subtokens (e.g., <_pal> <at> <able>) of a new token (e.g., <_palatable>) do not necessarily store the semantics of the specific concatenated form . Instead, their contexualized representation is gradually constructed in the Transformer layers via a neural detokenization process (Elhage et al., 2022; Lad et al., 2024; Kaplan et al., 2025) and then attended to by other token positions in the sequence, which are influenced by this contextualized representation. The various prior approaches that solely rely on information encoded in the embedding matrices of a model thus ignore the crucial functional knowledge stored in the attention and feed-forward weights of Transformer layers, which imposes a fundamental limitation on their performance.
@@ 3.2 Method Description
Let be the tokenizer of a pretrained Transformer. Given a new token string with original subtokens , we require a small corpus of example sequences each containing the new token string . We denote the new tokenizer, which includes , as , and the tokenization of using or as and , respectively. The target for learning the embedding of are the hidden states at a specified layer produced by the language model when processing , which we denote as . We wish to find an embedding for a new token such that hidden states produced by processing the input sequence instead of are similar to our target . We denote these new hidden states using our new embedding as . As the two input sequence indices are not aligned due to replacing subtokens by a single new token, we define a set of mapped positions , where and are the positions of the same token in and , respectively. Additionally, only includes pairs where position in would attend to . We use subscripts such as to signify the hidden state at the -th token position. We learn by minimizing the mean-squared error between hidden states for a given target layer :
@@ 3.3 Further Details
Like other methods optimizing embeddings based on contexts, our method also needs input sequences. Firstly, we note that randomly sampling texts from a domain-specific or general corpus is inefficient for our goal of learning newly added embeddings: typically, the new tokens will only make up a small fraction, so most gradient updates will actually not affect the new input embeddings at all and simply learn to minimize their log-probability in the case of output embeddings. As we aim to have a fast method, we need a better approach. We propose two different approaches: (1) Our main approach is to simply retrieve snippets that contain our target tokens from a domain-specific or general corpus. This can be implemented efficiently using the algorithm proposed by Aho & Corasick (1975). Then we can truncate the snippets to a small window around our target token to optimize computational efficiency. (2) For causal language models, we can simply generate snippets by prompting the model with our target token. We provide implementation details in Section B.7. In our main experiments, we focus on the first approach but note the availability of the second approach in case a reference corpus is not available. In most cases of domain and language adaptation, the availability of such a corpus is a reasonable assumption. Nevertheless, we study an ablation instead using the generative approach in Section 5.2.
@@
Since our method backpropagates gradients back from the hidden states, we do not learn output embeddings with our distillation-based objective. In fact, this is not possible, as our new tokens are not part of the original model that serves as the “teacher”. In practice, for learning output embeddings, we can simply add a next-token prediction objective just for the output embeddings at a minimal computational overhead or freely combine our method with any other method for initializing output embeddings. Recent work suggests that input and output embeddings should in fact be treated differently (Nakash et al., 2025; Huang et al., 2025). Indeed, the Transformer architecture can handle tokens that are input-only. Therefore, if not paired with an additional initialization method for output embeddings, we have the choice of either adding new tokens only to the input embedding matrix or – for compatibility with existing frameworks – setting their output embeddings to a vector of zeros.
@@
We employ a simple setup and use the AdamW (Kingma & Ba, 2017; Loshchilov & Hutter, 2019) optimizer for all trainable parameters. We set the batch size to optimize throughput and run all experiments on Nvidia H100 80GB GPUs. We do not use weight decay and maintain a constant learning rate with a linear warmup. For fair comparison, we sweep for the best learning rate for all methods that require a learning rate. Since our method is aimed to serve as an “initialization” rather than as full-scale further training, we restrict the number of example sequences to a maximum of 25 per target token and truncate to a context length of 50 tokens. These restrictions ensure that our method is quick to run, initializing 2,500 new tokens on a single GPU in under 10 minutes. We use the same data for all training-based methods, including baselines and variations of our method. We provide all details in Appendix B.
@@ 4.1 Evaluation
Common use cases for adding new tokens to a pretrained model’s tokenizer are language and domain adaptation, especially for the biomedical domain (Poerner et al., 2020; Gee et al., 2023; Hasan et al., 2024; Singh et al., 2024; Balde et al., 2024). This is because it is a particularly challenging domain with highly complex domain-specific terminology, which can serve as a benchmark to stress test embedding initialization methods. Therefore, we evaluate our method on a collection of standard benchmarks in the biomedical domain (Pal et al., 2024) and add frequently occurring words as new tokens. Additionally, for a more in-depth evaluation of the quality of new representations provided by our methods, we prompt the adapted models to generate definitions for such new tokens and evaluate their quality with an LLM-as-a-Judge (Li et al., 2023). We also evaluate Token Distillation for language adaptation, choosing French as an exemplary language. We evaluate our method on a set of multiple-choice benchmarks from FrenchBench (Faysse et al., 2025) and add frequently occurring words as new tokens, as in the biomedical domain adaptation experiments.
Furthermore, to push the limits of embedding initialization methods, we move beyond domain and language adaptation and apply our method to multi-word tokens, which have recently gained further interest (Gee et al., 2023; Liu et al., 2025; Huang et al., 2025). We follow the zero-shot tokenizer transfer setting in our evaluations (Minixhofer et al., 2024), i.e., we evaluate initialization methods without further training on a corpus, as we want to directly judge the quality of the resulting representations.
We select target tokens that actually occur frequently in the chosen benchmarks so that we can effectively judge the quality of newly generated embeddings. We select words that occur more frequently than a threshold and additionally ensure that all individual benchmarks are represented. We report the full methodology in Section B.1. For our multi-word token experiments, we prompt gpt-o4-mini-high to generate suitable candidates (full details in Section B.3). For our experiments in the biomedical domain, we retrieve contexts from ncbi/pubmed as a reference corpus. For our experiments on language adaptation to French, we use HuggingFaceFW/fineweb-2. For our multi-word token experiments, we prompt the original models to generate sequences containing the new tokens. For the LLM-as-a-Judge evaluations, we use Llama3.3-70B-Instruct as the judge model. We evaluate the general correctness of the generated definitions to test the quality and completeness of the resulting representations as well as semantic similarity with the target model’s output (Villegas et al., 2025), as inducing as little behavior change as possible can be an important desideratum.
@@ 4.2 Modeling
We conduct experiments using a wide range of open-weight model checkpoints to ensure our results are not merely a function of peculiarities in any specific model’s embedding space. In particular, we choose the following models: Mistral-7B-v0.1, Llama3-8B, Llama3-8B-Instruct, Llama3.1-8B, Llama3.1-8B-Instruct, Llama3.2-3B, Llama3.2-3B-Instruct, and OLMo2-7B-1124-Instruct. In the remainder of the paper, we denote “-Instruct” variants with a simple “-i” suffix. Through this extensive evaluation, we study different model families, model sizes (3B, 7B, and 8B), instruction-tuned and base models, as well as models with separate input/output embeddings and tied (shared) embeddings between input and output. For our experiments on French language adaptation, we run experiments with the Mistral-7B-v0.1 and Llama3-8B(-Instruct) models and additionally consider Qwen3-8B-Base, which has received more extensive multilingual training.
@@
For all results, we report the performance of the unmodified base model using the original tokenization. To establish a lower bound, we also report results for initializing the new token embeddings randomly from a normal distribution with a per-channel mean and standard deviation of the original embeddings (Hewitt, 2021). Furthermore, we report results for the commonly used method of taking the subtoken mean (Sachidananda et al., 2021; Koto et al., 2021; Gee et al., 2022) as initialization for a new token, which has been shown to perform similarly to more sophisticated initialization methods that also use a weighted average of existing embeddings (Minixhofer et al., 2024; Yamaguchi et al., 2024b).
Since our proposed method conducts a short optimization on a few reference sequences per new token, we compare against the common approach of training embeddings using causal language modeling using the same data. Specifically, we report results for “classic” embedding tuning using the next-token prediction objective – denoted as NTP (tune all embeddings) – as well as masking updates to the original embeddings such that only new embeddings are optimized (denoted as just NTP), which corresponds to the method used by Lampinen & McClelland (2018). These methods use the subtoken mean as their starting point. A strong alternative method for obtaining embeddings for new tokens are hyper-networks specifically pretrained for this task. We report results for ZeTT (Minixhofer et al., 2024) using their provided hyper-network checkpoints. We also compare against “Tokens to Words” (Kaplan et al., 2025), which relies on PatchScopes (Ghandeharioun et al., 2024) to extract hidden states and learns mappings from these to the embedding spaces. Unlike our Token Distillation approach, it hinges on a successful PatchScopes extraction of hidden states for each particular token. In our experiments, “Tokens to Words” could not be applied to 10–75% of tokens due failed PatchScope extractions, skewing the comparison with other methods that do not suffer from this limitation. We therefore report its results separately (Section A.1); Token Distillation performs better on average while also covering all added tokens. We run five random seeds for all methods that include randomness and report the mean and standard deviation.
@@
As the initial embedding input into our method, we also use the subtoken mean. We investigate alternatives to this choice in Section 5.2. Our proposed objective can also be combined with the NTP-based objectives. We report results for combining Token Distillation and NTP. Since the NTP and Token Distillation objectives have different scales, a simple addition is suboptimal. In fact, NTP is usually of larger magnitude, which leads it to overpower Token Distillation. Therefore, we also consider an “autoscaled” variant Token Distillation + NTP where the NTP loss is scaled by = stop_gradient(loss_awedist / loss_ntp) before summing. Here, stop_gradient prevents gradient flow through the scaling factor .333Multi-objective optimization is a rich field with extensive literature, we leave a thorough investigation of alternatives such as GradNorm (Chen et al., 2018) for future work.
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP (tune all embeddings) 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 NTP (Lampinen et al., 2018) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
Method Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Original tokenization 99.8 96.5 100.0 94.3 100.0 98.4 100.0 94.7 100.0 98.4 100.0 93.2 100.0 93.8 99.4 96.9 99.9 95.8 Random (Hewitt, 2021) 0.0 0.4 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 Mean (Gee et al., 2022) 2.5 4.3 16.8 16.2 17.6 20.7 25.6 25.6 23.2 28.3 15.2 15.2 12.1 14.6 20.1 24.2 16.6 18.6 NTP (tune all embeddings) 26.8 30.3 33.2 39.5 34.4 43.4 28.5 35.0 41.4 50.0 19.7 20.7 18.6 21.1 22.5 26.6 28.1 33.3 NTP (Lampinen et al., 2018) 49.8 58.2 45.3 49.6 48.4 58.0 58.6 65.6 60.2 69.9 51.6 58.4 50.0 58.8 52.1 57.0 52.0 59.4 ZeTT (Minixhofer et al., 2024) 63.5 68.6 69.5 73.6 70.7 80.3 – – – – – – – – – – – – Token Distillation 79.7 85.2 72.7 76.2 79.7 91.0 75.8 80.1 81.6 89.8 0.0 0.0 75.0 83.0 83.8 89.5 68.5 74.4 Token Distillation + NTP 77.3 82.6 68.4 73.8 75.8 86.3 72.3 79.9 77.3 86.1 62.3 69.1 65.2 75.8 65.4 74.6 70.5 78.5 Token Distillation + NTP 76.8 82.0 71.7 75.2 80.9 91.6 77.5 81.2 81.1 89.5 72.7 79.5 71.9 80.5 81.2 86.9 76.7 83.3
Mistral-7B Llama3-8B Llama3-8B-i Qwen3-8B Avg. Original tokenization 69.5 69.4 72.1 81.7 73.2 Random (Hewitt, 2021) 51.0 46.1 51.5 62.1 52.7 Mean (Gee et al., 2022) 56.3 58.4 61.7 69.6 61.5 NTP (Lampinen et al., 2018) 64.7 67.0 70.1 81.2 70.8 ZeTT (Minixhofer et al., 2024) 66.1 61.3† 65.2† – – Token Distillation 68.5 68.9 72.9 81.5 72.9